OpenStudy (anonymous):

A square pyramid has a base side of 15 in. and a slant height of 16 in. The lateral area of the pyramid is _____ in.². 60 in.² 120 in.² 240 in.² 480 in.²

OpenStudy (anonymous):

@brusmack

OpenStudy (anonymous):

Hello? pooja195?

pooja195 (pooja195):

LA = ½ps where p is the perimeter of base. and s is the Slant height.

pooja195 (pooja195):

Perimeter = 15 + 15 + 15 + 15 = 60 LA = ½(60)(16)

pooja195 (pooja195):

Option D

pooja195 (pooja195):

yr answer

OpenStudy (anonymous):

kk thank you :) i get it now :)

OpenStudy (anonymous):

480 is the answer :)

OpenStudy (anonymous):

A triangular pyramid has a volume of 1 ft³. If all its dimensions are doubled, the volume of the new pyramid will be _____ ft³. 6 8 16 24

OpenStudy (anonymous):

can you tell me if the answer is 8 im just checking

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

tell me how u go yr answer

OpenStudy (anonymous):

i just timed 1 * 8 im confused now

OpenStudy (anonymous):

well 3.14

pooja195 (pooja195):

alright hol on

OpenStudy (anonymous):

okay

pooja195 (pooja195):

Generally, finding volume involves multiplying 3 dimensions together. You're doubling all these dimensions. So, instead of a*b*c you get (2a)*(2b)*(2c).

OpenStudy (anonymous):

alright so now what

OpenStudy (anonymous):

so the answer is 6

pooja195 (pooja195):

so yr answer would be corrrrrect

OpenStudy (anonymous):

oh hah i jsut made a mistake yes thank you

OpenStudy (anonymous):

A cone has a base area of 3 in² and a lateral area which is 1.5 times bigger than the base area. The surface area of the cone is _____ in². 4.5 6.0 7.5 9.0

OpenStudy (anonymous):

how bout this one

OpenStudy (anonymous):

is 4.5 correct i just timed them

pooja195 (pooja195):

Lateral area = 1.5 (3) = 4.5 Surface area = lateral area + base area = 4.5 + PI r^2 = 4.5 + 3 = 4.5+3 = 7.5

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

ah okay

OpenStudy (anonymous):

What is the lateral area of a cylinder with a radius of 10 in and a height of 6 in? Use 3.14 for

OpenStudy (anonymous):

60 in² 376.8 in² 1004.8 in² 628 in²

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

if u post the question together it would help :-)

pooja195 (pooja195):

just a thought

OpenStudy (anonymous):

oh sorry. well i jsur seperated the answers from the question

OpenStudy (anonymous):

so what did u get

pooja195 (pooja195):

By 'lateral area' I assume you mean the curved area. Lateral area = height * circumference = 6 in * 2 * 3.14 * 10 in = 376.8 in^2

OpenStudy (anonymous):

Thank you ..

OpenStudy (anonymous):

i need help with one more

OpenStudy (anonymous):

Find the area of the following figure.